b1az3:

What is the value of x in the following equation? 8 -8 -6 6

b1az3:

1 attachment
AZ:

The progress we've made so far: https://questioncove.com/study#/updates/6048d4d93e1d45f60f901600 We determined that the left side was \(\Large 3^{16x}\) but what is the right side simplified to?

AZ:

let me give you colors again \( \Large \dfrac{a^{\color{green}{b}}}{a^{\color{orange}{c}}} = a^{\color{green}{b}-\color{orange}{c}} \)

AZ:

Here's an example \(\Large \dfrac{2^{12x}} {2^{8}} = 2^{(12x - 8)}\) can you now do it for YOUR question?

AZ:

so then for YOUR question, what is \(\Large \dfrac{3^{20x}}{a^{24}} = ?\) just subtract the exponents and write it like I did in that example I just gave you